### v1

@@ 3.2. Design Approaches for Drop’n’Collect
Locative interactions, such as Pokémon Go [^1] https://pokemongo.com/, a mobile game, pin interactive elements to specific geolocations and require users to physically go there. And locative narratives are those leveraging locative interactions for content consumption and creation (nisiAuthoringLocativeNarratives2022). As mentioned in section 2.2, they can offer situated content consumption and an engaging and entertaining user experience through physically engaged interactions. It is also adopted by commercial platforms such as AllTrails [^2] https://www.alltrails.com/, a hiking experience sharing and trail authoring platform., as well as HCI works for history preservation (cheverstSupportingConsumptionCoauthoring2017), interactive fiction (millardTireeTalesCooperative2017; millardCanyonsDeltasPlains2013), random short messages (liuPinsightNovelWay2018; cranshawJourneysNotesDesigning2016; juBreadcrumbSNSAsynchronous2015), and specialized information (changTourgetherExploringTourists2020; siriarayaPalmHappinessLocationBased2022). Thus, we anticipate that locative narratives achieve a similar effect of meaningful content sharing in the context of collective memory co-contribution (DR5). In addition, memories are often strongly related to specific locations (bennettTopoTilesStorytellingCare2015; tuanTopophiliaStudyEnvironmental1990). Being situated in the location can trigger people’s recalling of memories and the accompanying emotions (tuanTopophiliaStudyEnvironmental1990) (DR1). Lastly, location information is an essential feature of collective memory that holds “environments” and “meanings” beyond mere space (lakCollectiveMemoryUrban2019). It can serve as a link across spatially related memory pieces and the bearer of people’s feelings and thoughts (tuanTopophiliaStudyEnvironmental1990) (DR4).
@@
Skeuomorphism refers to the UI design technique of imitating similar artifacts in another material (pageSkeuomorphismFlatDesign2014). It is frequently used in HCI designs for personal reflection (sunPostcardYourFood2020; gerritsenMailingArchivedEmails2016), memory anchors (shojiMuseumExperienceSouvenir2024; kangTieMemoriesEsouvenirs2022), and so on. For example, the postcard is an effective skeuomorphic object to facilitate the reflection of the past (sunPostcardYourFood2020; gerritsenMailingArchivedEmails2016). Encompassing the abstract concept of memory in familiar and everyday skeuomorphic concepts not only allows engaging interactions (DR5) but also stresses the meaningfulness of their sharing (DR1).
@@
Finally, we borrowed the construct “spread of activation” from psychology to describe the mechanism of inciting and discovering more associated memory content within the community. This was intended to complement locative narratives that center primarily on memories about a particular spatial region. As a cue to direct the users (i.e., how the “network” is connected in the analogy, DR4), we leveraged Lak et al.’e framework of collective memory’s core elements, such as locations’ meanings, things people do, personal contexts, and history over time, as they are the natural cues that associate individual memories in a community (lakCollectiveMemoryUrban2019). Analogous to “activating” a selection of “source nodes” and propagating the “activation” to other “nodes in a network” (crestaniSpreadActivationTheory2012), we wanted to engage them in exploring other related memories that were spatially and temporally distributed, and raise their awareness of the multi-facet nature of collective memory. Consequently, it could enhance the coverage and diversity of their conceptualization of the collective memory and their contribution to it (DR2).
@@ 3.3. Features and Implementation
[t]0.4 Refer to caption : {subcaptionblock}[t]0.5 Refer to caption : {subcaptionblock}[T]
Figure 5. Browsing memories in both systems:
@@
{subcaptionblock}
[T]0.58 Refer to caption Refer to caption Refer to caption : {subcaptionblock}[T]0.4 Refer to caption :
@@ 3.3.1. Memory Exploration and Creation
In Drop’n’Collect, users’ memory posts are tied to their current physical geolocations, and they can only see the memory posts nearby. This way, the exploration and curation of memories occurs in conjunction with physical exploration of the campus space, forming a locative narrative experience (nisiAuthoringLocativeNarratives2022). We intend to leverage the in-person witness of the physical surroundings to stimulate writers’ recollection and augment readers’ impression (DR1). And we also intend to encourage the process of physical exploration, which may add value to the memories users discover and read (DR1, DR5)
@@ 3.3.2. Visual Elements
We propose two concrete skeuomorphic metaphors for Drop’n’Collect: postcards and memory boxes. Postcards are often used to hold text expressions and graphics about a particular experience (gerritsenMailingArchivedEmails2016) for meaningful reflection (sunPostcardYourFood2020) and sharing with others (kellyUnderstandingParticipationOpportunities2012). Each memory post is made into the shape of a postcard, with a required photograph as the cover, the memory’s geolocation as the address, and the author’s avatar as the stamp (Figure 10). Viewers can flip between the photo and content sides of the postcard.
Memory boxes afford to contain and transmit memory-related inventories (aaliMemoryBoxesExperimental2014) (DR4), and they afford the interactions of opening and closing (kleinbergerSupportingElderConnectedness2019), which may be integrated into our system to increase interactivity and engagement (DR5). The memory box is designed to be an interactive exercise of “opening a post”: users tilt their mobile phones upward to mimic the action of opening a box. We postulate that the tilting-up action may also guide the user to lift up the head, bringing the vision from the screen to the physical surroundings (DR4). To link this metaphor to the postcard, the memory box contains the “spirit” of a memory that refreshes the time-washed photograph back to normal. Similarly, when users post a new memory, a reversed animation stores the “spirit” into the memory box. A thank-you message also signifies the user’s contribution to the community’s collective memory (DR3).
@@ 3.3.4. Memory Metadata
To complement the core elements of collective memory-related content (see section 3.2), we cater to the activities and personal contexts as posts’ metadata in Drop’n’Collect. Two specific aspects of personal contexts, emotions and weather, were chosen due to their usual association with the memory experience (sasDesigningCollectiveRemembering2006). The activity is free-form text, whereas the aspects of the personal context are in options. The emotions are from the basic emotion theory (anEmoWearExploringEmotional2024) (happy, angry, afraid, surprised, sad, calm, and neutral), and the weathers are from Apple’s list of weather icons [^4] https://support.apple.com/guide/weather-mac/symbols-used-in-weather-apdw00197118/mac, excluding mixed weathers and those not available locally (clear, cloudy, rainy, windy, foggy, not applicable). To further lower their cognitive load of additional input (DR3), we wrote simple GenAI prompts to pre-fill them based on the post content.
@@
In the forum system, since we do not explicitly practice Lak et al.’s framework about collective memory’s core elements (lakCollectiveMemoryUrban2019), we do not offer metadata in specific dimensions. To provide a matching feature for the sake of informativeness, users can write free-form text as hashtags — similar to many real-world forum designs — beneath the post. It does not specifically recommend far-away posts, as users always browse all posts without geolocation limits. We expect that the metadata in the forum system may add less value compared to Drop’n’Collect, because users may have already written more informative content, and they are already accessible to all the posts.
@@ 4.1. Scenario
Among existing literature, we have identified two primary scenarios when investigating a community’s collective memory. Some research focuses on the scenario of users’ explicit and purposeful acts that contribute to collective memory, as well as the possible use of various intentional designs for collective memory development (jonesCoconstructingFamilyMemory2018; lindleyNarrativeMemoryPractice2009; sasDesigningCollectiveRemembering2006; jonesDesigningSocialMemory2016). Meanwhile, many other studies adopt the case where collective memory is an organic by-product on general-purpose platforms (e.g., Wikipedia (ferronCollectiveMemoryBuilding2011; twymanBlackLivesMatter2017), social media (aldamenStimulationCollectiveMemory2024; zhangCrisisCollectiveMemory2020; khlevnyukNarrowcastingCollectiveMemory2019)), discussing its possible integration into people’s generic online social activities.
@@ 4.2. Methods
The interview transcripts underwent a thematic analysis. We started deductively with three high-level themes of user interactions relevant to our research questions and interview structure (braunThematicAnalysisPractical2022): participants’ contribution of memory content (RQ1, RQ2), their reading and comprehension of others’ content (RQ1, RQ2), and feature usability (RQ2, RQ3). Guided by the three themes, we then inductively extracted codes to capture the emerging insights (braunThematicAnalysisPractical2022), such as the characteristics of user posts, their interactional patterns, and users’ mental interpretations. During the coding, two researchers independently coded the data first. Two subsequent rounds of joint discussions resolved the conflicts and confusions between similar codes across themes. The final codes are presented in a thematic map (ahmedUsingThematicAnalysis2025) (Figure 15), and elaborated in the following section.
The questionnaires were to supplement our understanding of the user feedback. In line with the major benefit of collective memory, we inspected whether the systems could facilitate connectedness with community members (sasDesigningCollectiveRemembering2006) with the Campus Connected Scale (CCS) (leeCulturalOrientationMulticultural2000) and the collective identity (roedigerCollectiveMemoryNew2015; eyermanPresentCultureTransmission2004) with Collective Self-Esteem Scale (CSES) (luhtanenCollectiveSelfEsteemScale1992). (We excluded the “membership” and “public” sections of CSES. The first overlapped with CCS. The second is about the public image of being part of the group, and was out of this study’s scope.) They were sampled before and after the experience period, and we calculated the difference between the two samples. In addition, we used UEQ-S (schreppDesignEvaluationShort2017) plus feature-specific customized questions only in the exit questionnaire to see how the current design supported the practical co-contribution of collective memory.
We chose the Mann-Whitney U test (mannTestWhetherOne1947) to compare the questionnaire data (i.e., the increments in CCS and CSES items, the direct responses to other items) for the following reasons. We first used Shapiro-Wilk normality tests (ghasemiNormalityTestsStatistical2012) to verify that either group’s responses to most questions did not follow normal distributions. (Only two out of 44 questions had both group’s responses in normal distributions, so we still considered the overall data not normal.) This combined with the small sample size (below 50) suggested nonparametric tests (krzywinskiNonparametricTests2014; kapteinPowerfulConsistentAnalysis2010). Furthermore, different individuals divided in two groups form independent samples, for which the Mann-Whitney U test is the most common nonparametric test method (nacharMannWhitneyTestAssessing2008). Prior HCI scholars also validate and recommend the Mann-Whitney U test for Likert scale data in common HCI studies (kapteinPowerfulConsistentAnalysis2010).
@@ 4.3. Participants
While we were sampling and grouping participants, we ensured that each group had both newcomers to the university and experienced residents, so that they were more likely to have both memory sharing and discovery. Thus, we added a self-reported “year (of residence) on campus” metric in the recruitment survey. As we grouped the participants, we ensured that this attribute followed similar distributions in both groups (shown in Table 2). Group A has a mean of 3.16 and std of 2.58, and Group B has a mean of 2.79 and std of 2.44. A further Mann-Whitney U test on this attribute also verified statistical insignificance ($p=0.801$). (Similar to section 4.2, we used the Shapiro-Wilk test (ghasemiNormalityTestsStatistical2012) to verify non-normality and followed the same subsequent rationale to choose the Mann-Whitney U test.)
@@ 4.4. Procedure
Before the evaluation, we inserted the same 23 memories as seed content to bootstrap engagement at the beginning of the evaluation (cranshawJourneysNotesDesigning2016). They were authentic memories of three researchers and three external authors recruited via on-campus advertisement. We deliberately recorded memories of different places on campus and balanced text and audio posts. The same seed content was stored in each system’s database, and the audio posts were transcribed into text in the forum system.
@@ 5. Results
*: These questions were negated in our survey to stay consistent with the original literature (leeCulturalOrientationMulticultural2000; luhtanenCollectiveSelfEsteemScale1992). For the sake of presentational clarity, we reverse the question text and mirror the ratings in this table, so that larger numbers are always better.
@@ 5.1. Participants’ Contribution of Memory Posts
We first looked into how users posted memory posts in both systems, which most directly reflected their conception of co-contributing collective memory. At first sight, Drop’n’Collect users created significantly more posts that denoted the “presence”—anything that users immediately saw, did, or thought of. Consequently, there were more short contents like “Late night working [with a photo of a desk]”, “Rainy~ [with a photo of an open-air plaza]”, and “go swimming [with a photo of the way to the swimming pool]”. Interestingly, A8 saw a basketball with headphones attached to it on a nearby desk, and posted a picture of it with the caption “What a handsome basketball!” On the contrary, the forum posts were generally longer, more topically diverse, and there were more personal recollections related to the locations. For example, B7 wrote “Staring at the horizon from the library. Is the breeze more bitter than the schoolwork? The scenery was the reason I chose this university without hesitation. Will I love the scenery as now after several years?” According to our observations, these memory outputs in the forum system present similar styles to those reported by other studies of collective memory on general-purpose social platforms: detailed, contextual, sometimes emotionally charged or polemical (twymanBlackLivesMatter2017; glawionRememberingWorldWar2023; zhangCrisisCollectiveMemory2020).
@@
[t]0.32 Refer to caption {subcaptionblock}[t]0.32 Refer to caption {subcaptionblock}[t]0.32 Refer to caption :

### v2

@@ 3.2. Design Approaches for Drop’n’Collect
Locative interactions, such as Pokémon Go [^1] https://pokemongo.com/, a mobile game, pin interactive elements to specific geolocations and require users to physically go there. And locative narratives are those leveraging locative interactions for content consumption and creation (Nisi, 2022). As mentioned in section 2.2, they can offer situated content consumption and an engaging and entertaining user experience through physically engaged interactions. It is also adopted by commercial platforms such as AllTrails [^2] https://www.alltrails.com/, a hiking experience sharing and trail authoring platform., as well as HCI works for history preservation (Cheverst et al., 2017), interactive fiction (Millard and Hargood, 2017; Millard et al., 2013), random short messages (Liu et al., 2018; Cranshaw et al., 2016; Ju et al., 2015), and specialized information (Chang et al., 2020; Siriaraya et al., 2022). Thus, we anticipate that locative narratives achieve a similar effect of meaningful content sharing in the context of collective memory co-contribution (DR5). In addition, memories are often strongly related to specific locations (Bennett et al., 2015; Tuan, 1990). Being situated in the location can trigger people’s recalling of memories and the accompanying emotions (Tuan, 1990) (DR1). Lastly, location information is an essential feature of collective memory that holds “environments” and “meanings” beyond mere space (Lak and Hakimian, 2019). It can serve as a link across spatially related memory pieces and the bearer of people’s feelings and thoughts (Tuan, 1990) (DR4).
@@
Skeuomorphism refers to the UI design technique of imitating similar artifacts in another material (Page, 2014). It is frequently used in HCI designs for personal reflection (Sun et al., 2020; Gerritsen et al., 2016), memory anchors (Shoji et al., 2024; Kang et al., 2022), and so on. For example, the postcard is an effective skeuomorphic object to facilitate the reflection of the past (Sun et al., 2020; Gerritsen et al., 2016). Encompassing the abstract concept of memory in familiar and everyday skeuomorphic concepts not only allows engaging interactions (DR5) but also stresses the meaningfulness of their sharing (DR1).
@@
Finally, we borrowed the construct “spread of activation” from psychology to describe the mechanism of inciting and discovering more associated memory content within the community. This was intended to complement locative narratives that center primarily on memories about a particular spatial region. As a cue to direct the users (i.e., how the “network” is connected in the analogy, DR4), we leveraged Lak et al.’e framework of collective memory’s core elements, such as locations’ meanings, things people do, personal contexts, and history over time, as they are the natural cues that associate individual memories in a community (Lak and Hakimian, 2019). Analogous to “activating” a selection of “source nodes” and propagating the “activation” to other “nodes in a network” (Crestani, 2012), we wanted to engage them in exploring other related memories that were spatially and temporally distributed, and raise their awareness of the multi-facet nature of collective memory. Consequently, it could enhance the coverage and diversity of their conceptualization of the collective memory and their contribution to it (DR2).
@@ 3.3. Features and Implementation
[t]0.4 Refer to caption The left screenshot shows the map view of Drop’n’Collect with pins of nearby posts’ locations and a horizontal list of their photos. The right screenshot shows the list view of the forum with every post’s photo and content preview. {subcaptionblock}[t]0.5 Refer to caption Screenshots when viewing a memory post on two systems. For Drop’n’Collect, a postcard on the top is flippable between the photo side and the content side. Below is a horizontal list of check-in photographs, each being a user’s picture and an overlay of a fake hand holding the postcard. Further below is a horizontal list of sepia photos as the recommended memories. For the forum, vertically laid out are the text content, the photo, the hashtag list, the “like” button, the date and time of the post, and the author’s ID. {subcaptionblock}[T]
Figure 5. Browsing memories in both systems
@@
On the left are four screenshots showing the procedure of publishing a memory postcard in Drop’n’Collect: choose a photo, write or speak content on a virtual postcard, select emotions and weather, and optionally input an activity, confirm the location, and optionally clarify the location, see the thank-you message. On the right is one screenshot of publishing a memory post in the forum: write the text content, optionally add a photo, and optionally add hashtags. {subcaptionblock}
[T]0.58 Refer to caption Refer to caption Refer to caption Three screenshots that show the following animation: a green box opens, a yellow-white sparkling item appears from inside the box and flies to a sepia photograph, and the sepia photograph then recovers its original color. {subcaptionblock}[T]0.4 Refer to caption A camera view with the overlay at the bottom center of a fake hand holding a postcard.
@@ 3.3.1. Memory Exploration and Creation
In Drop’n’Collect, users’ memory posts are tied to their current physical geolocations, and they can only see the memory posts nearby. This way, the exploration and curation of memories occurs in conjunction with physical exploration of the campus space, forming a locative narrative experience (Nisi, 2022). We intend to leverage the in-person witness of the physical surroundings to stimulate writers’ recollection and augment readers’ impression (DR1). And we also intend to encourage the process of physical exploration, which may add value to the memories users discover and read (DR1, DR5)
@@ 3.3.2. Visual Elements
We propose two concrete skeuomorphic metaphors for Drop’n’Collect: postcards and memory boxes. Postcards are often used to hold text expressions and graphics about a particular experience (Gerritsen et al., 2016) for meaningful reflection (Sun et al., 2020) and sharing with others (Kelly and Gooch, 2012). Each memory post is made into the shape of a postcard, with a required photograph as the cover, the memory’s geolocation as the address, and the author’s avatar as the stamp (Figure 10). Viewers can flip between the photo and content sides of the postcard.
Memory boxes afford to contain and transmit memory-related inventories (Aali and Rogge, 2014) (DR4), and they afford the interactions of opening and closing (Kleinberger et al., 2019), which may be integrated into our system to increase interactivity and engagement (DR5). The memory box is designed to be an interactive exercise of “opening a post”: users tilt their mobile phones upward to mimic the action of opening a box. We postulate that the tilting-up action may also guide the user to lift up the head, bringing the vision from the screen to the physical surroundings (DR4). To link this metaphor to the postcard, the memory box contains the “spirit” of a memory that refreshes the time-washed photograph back to normal. Similarly, when users post a new memory, a reversed animation stores the “spirit” into the memory box. A thank-you message also signifies the user’s contribution to the community’s collective memory (DR3).
@@ 3.3.4. Memory Metadata
To complement the core elements of collective memory-related content (see section 3.2), we cater to the activities and personal contexts as posts’ metadata in Drop’n’Collect. Two specific aspects of personal contexts, emotions and weather, were chosen due to their usual association with the memory experience (Sas and Dix, 2006). The activity is free-form text, whereas the aspects of the personal context are in options. The emotions are from the basic emotion theory (An et al., 2024) (happy, angry, afraid, surprised, sad, calm, and neutral), and the weathers are from Apple’s list of weather icons [^4] https://support.apple.com/guide/weather-mac/symbols-used-in-weather-apdw00197118/mac, excluding mixed weathers and those not available locally (clear, cloudy, rainy, windy, foggy, not applicable). To further lower their cognitive load of additional input (DR3), we wrote simple GenAI prompts to pre-fill them based on the post content.
@@
In the forum system, since we do not explicitly practice Lak et al.’s framework about collective memory’s core elements (Lak and Hakimian, 2019), we do not offer metadata in specific dimensions. To provide a matching feature for the sake of informativeness, users can write free-form text as hashtags — similar to many real-world forum designs — beneath the post. It does not specifically recommend far-away posts, as users always browse all posts without geolocation limits. We expect that the metadata in the forum system may add less value compared to Drop’n’Collect, because users may have already written more informative content, and they are already accessible to all the posts.
@@ 4.1. Scenario
Among existing literature, we have identified two primary scenarios when investigating a community’s collective memory. Some research focuses on the scenario of users’ explicit and purposeful acts that contribute to collective memory, as well as the possible use of various intentional designs for collective memory development (Jones and Ackerman, 2018; Lindley et al., 2009; Sas and Dix, 2006; Jones, 2016). Meanwhile, many other studies adopt the case where collective memory is an organic by-product on general-purpose platforms (e.g., Wikipedia (Ferron and Massa, 2011; Twyman et al., 2017), social media (Aldamen and Abdul Jaleel, 2024; Zhang et al., 2020; Khlevnyuk, 2019)), discussing its possible integration into people’s generic online social activities.
@@ 4.2. Methods
The interview transcripts underwent a thematic analysis. We started deductively with three high-level themes of user interactions relevant to our research questions and interview structure (Braun and Clarke, 2022): participants’ contribution of memory content (RQ1, RQ2), their reading and comprehension of others’ content (RQ1, RQ2), and feature usability (RQ2, RQ3). Guided by the three themes, we then inductively extracted codes to capture the emerging insights (Braun and Clarke, 2022), such as the characteristics of user posts, their interactional patterns, and users’ mental interpretations. During the coding, two researchers independently coded the data first. Two subsequent rounds of joint discussions resolved the conflicts and confusions between similar codes across themes. The final codes are presented in a thematic map (Ahmed et al., 2025) (Figure 15), and elaborated in the following section.
The questionnaires were to supplement our understanding of the user feedback. In line with the major benefit of collective memory, we inspected whether the systems could facilitate connectedness with community members (Sas and Dix, 2006) with the Campus Connected Scale (CCS) (Lee and Davis III, 2000) and the collective identity (Roediger and Abel, 2015; Eyerman, 2004) with Collective Self-Esteem Scale (CSES) (Luhtanen and Crocker, 1992). (We excluded the “membership” and “public” sections of CSES. The first overlapped with CCS. The second is about the public image of being part of the group, and was out of this study’s scope.) They were sampled before and after the experience period, and we calculated the difference between the two samples. In addition, we used UEQ-S (Schrepp et al., 2017) plus feature-specific customized questions only in the exit questionnaire to see how the current design supported the practical co-contribution of collective memory.
We chose the Mann-Whitney U test (Mann and Whitney, 1947) to compare the questionnaire data (i.e., the increments in CCS and CSES items, the direct responses to other items) for the following reasons. We first used Shapiro-Wilk normality tests (Ghasemi and Zahediasl, 2012) to verify that either group’s responses to most questions did not follow normal distributions. (Only two out of 44 questions had both group’s responses in normal distributions, so we still considered the overall data not normal.) This combined with the small sample size (below 50) suggested nonparametric tests (Krzywinski and Altman, 2014; Kaptein et al., 2010). Furthermore, different individuals divided in two groups form independent samples, for which the Mann-Whitney U test is the most common nonparametric test method (Nachar, 2008). Prior HCI scholars also validate and recommend the Mann-Whitney U test for Likert scale data in common HCI studies (Kaptein et al., 2010).
@@ 4.3. Participants
While we were sampling and grouping participants, we ensured that each group had both newcomers to the university and experienced residents, so that they were more likely to have both memory sharing and discovery. Thus, we added a self-reported “year (of residence) on campus” metric in the recruitment survey. As we grouped the participants, we ensured that this attribute followed similar distributions in both groups (shown in Table 2). Group A has a mean of 3.16 and std of 2.58, and Group B has a mean of 2.79 and std of 2.44. A further Mann-Whitney U test on this attribute also verified statistical insignificance ($p=0.801$). (Similar to section 4.2, we used the Shapiro-Wilk test (Ghasemi and Zahediasl, 2012) to verify non-normality and followed the same subsequent rationale to choose the Mann-Whitney U test.)
@@ 4.4. Procedure
Before the evaluation, we inserted the same 23 memories as seed content to bootstrap engagement at the beginning of the evaluation (Cranshaw et al., 2016). They were authentic memories of three researchers and three external authors recruited via on-campus advertisement. We deliberately recorded memories of different places on campus and balanced text and audio posts. The same seed content was stored in each system’s database, and the audio posts were transcribed into text in the forum system.
@@ 5. Results
*: These questions were negated in our survey to stay consistent with the original literature (Lee and Davis III, 2000; Luhtanen and Crocker, 1992). For the sake of presentational clarity, we reverse the question text and mirror the ratings in this table, so that larger numbers are always better.
@@ 5.1. Participants’ Contribution of Memory Posts
We first looked into how users posted memory posts in both systems, which most directly reflected their conception of co-contributing collective memory. At first sight, Drop’n’Collect users created significantly more posts that denoted the “presence”—anything that users immediately saw, did, or thought of. Consequently, there were more short contents like “Late night working [with a photo of a desk]”, “Rainy~ [with a photo of an open-air plaza]”, and “go swimming [with a photo of the way to the swimming pool]”. Interestingly, A8 saw a basketball with headphones attached to it on a nearby desk, and posted a picture of it with the caption “What a handsome basketball!” On the contrary, the forum posts were generally longer, more topically diverse, and there were more personal recollections related to the locations. For example, B7 wrote “Staring at the horizon from the library. Is the breeze more bitter than the schoolwork? The scenery was the reason I chose this university without hesitation. Will I love the scenery as now after several years?” According to our observations, these memory outputs in the forum system present similar styles to those reported by other studies of collective memory on general-purpose social platforms: detailed, contextual, sometimes emotionally charged or polemical (Twyman et al., 2017; Glawion, 2023; Zhang et al., 2020).
@@
[t]0.32 Refer to caption {subcaptionblock}[t]0.32 Refer to caption {subcaptionblock}[t]0.32 Refer to caption (a) The forum has considerably more posts mentioning the physical surroundings and linking to the past. Drop’n’Collect posts have slightly more mentions in other aspects. (b) Drop’n’Collect posts are slightly inclined to mention no more than two aspects compared to Forum. (c) Drop’n’Collect has considerably more posts about food and schoolwork, and the forum has considerably more posts about recollection and information inquiry. The forum has slightly more posts in all other topics than Drop’n’Collect.